OpenStudy (quickstudent):

Is this correct? (x^5 - x^4 + 3x^3 + x^2) - (9x^5 - 4x^4 + 3x^3 - 5x) x^5 - x^4 + 3x^3 + x^2 - 9x^5 + 4x^4 - 3x^3 + 5x x^5 - 9x^5 - x^4 + 4x^4 + 3x^3 - 3x^3 + x^2 + 5x -8x^5 + 3x^4 + x^2 + 5x

OpenStudy (quickstudent):

@mathmate

OpenStudy (mathmate):

I suppose you mean: (x^5 - x^4 + 3x^3 + x^2) - (9x^5 - 4x^4 + 3x^3 - 5x) =x^5 - x^4 + 3x^3 + x^2 - 9x^5 + 4x^4 - 3x^3 + 5x =x^5 - 9x^5 - x^4 + 4x^4 + 3x^3 - 3x^3 + x^2 + 5x =-8x^5 + 3x^4 + x^2 + 5x Yes, this is correct! Well done!

OpenStudy (quickstudent):

Thanks:)

OpenStudy (mathmate):

You're welcome! :)